OpenStudy (anonymous):

Type the formula for the number of permutations of n things taken r at a time. p(n,r)=

OpenStudy (anonymous):

2x^2+4=5-2x^2

OpenStudy (anonymous):

\[ \frac {n!}{(n-r)!} \]

Parth (parthkohli):

\( \color{Black}{\Rightarrow P(n,r) = \Large {n! \over (n - r)!} }\) lol had to give the answer out here.

Parth (parthkohli):

Forgive me, Mr. Fool